# Bulk Motions in the Black Hole Jet Sheath as a Candidate for the Comptonizing Corona

### v1

@@ 3 GRRMHD simulation results
Figure 3: Realizations of plasmoid-mediated reconnection layers identified in the jet sheath by the conditions $\sigma_{\rm h}>0.15$ and $\beta>0.1$ at various times (in different rows). Panels [a-d] show $\sigma_{\rm h}$; panels [e-h] show $\beta$; panels [i-l] show the current density. The bulk energy per particle in this region (i.e., where $\sigma_{\rm h}>0.15$ and $\beta>0.1$) is shown in panels [m-p]. In all the panels, the black hole is centered at $[x,z]=[0,0]$, the red dashed curve is the ergosphere, and the red solid curve is the inner light surface. Click here for a YouTube video of some of these quantities.
@@ 3.1 General properties of the jet sheath
In the top panels of Fig. 2, we present two-dimensional snapshots of the plasma density, the dimensionless temperature ($p/\rho c^{2}$), and the bulk energy per unit rest mass energy ($\Gamma-1$) at time $Tc/R_{\rm g}=2870$. One can identify the jet sheath as the region where the dimensionless temperature is $p/\rho c^{2}\sim 0.1$ and the bulk energy is $\Gamma-1\sim 1$. In the jet sheath, the flow transitions from being ultra-relativistically fast and hot in the jet to non-relativistic and cold in the disk wind. The plasma mass density is $\rho\sim 0.1$ (in code units) at the jet-disk interface. In the following sections, we present a different criterion—based on $\sigma_{\rm h}$ and $\beta$—to identify the regions with reconnection-like features in the jet sheath.
@@ 3.2.1 Location and geometry
Note that the presence of a current sheet does not guarantee fast dissipation. However, a current sheet showing plasmoid chains is undergoing active reconnection, which implies fast dissipation. Most of the time, long current sheets in the jet sheath are seen to harbor chains of plasmoids, which appear as prominent regions with $\beta\gtrsim 1$. Most of the plasmoids residing in/near the sheath at a given time were born in the sheath itself. However, there are brief moments (of duration $\sim 10\,R_{\rm g}/c$) when plasmoids are not formed in situ in the sheath: these are the times immediately following a flux eruption. During those moments, the region adjacent to the sheath (towards the disk, with $\sigma_{\rm h}<0.15$) may still harbor large plasmoids ($\sim 5\,R_{\rm g}/c$), which were created at the base of the jet during the prior flux eruption. These plasmoids can grow to large sizes (${\cal O}(10)\,R_{\rm g}$) while propagating outwards. Some examples appear in the top panel of Fig. 2 as the closed loops of magnetic field lines (e.g., at $[x/R_{\rm g},z/R_{\rm g}]=[16,7],[24,22]$). Numerical simulations have revealed that some of these large plasmoids move along a spiral trajectory around the jet (Nathanail et al., 2020; Ripperda et al., 2022; El Mellah et al., 2023).
Given the copious presence of plasmoids and reconnecting layers in the jet sheath, we regard this interface as an important site of energy dissipation. As such, it may be a potential candidate for the Comptonizing corona. More quantitatively, we identify the reconnection-powered dissipation regions in the jet sheath by the following combination: plasma beta $\beta>1$ and hot magnetization $\sigma_{\rm h}>0.15$. The former threshold excludes the jet core (which has lower plasma beta, and is unphysically affected by numerical density and pressure floors), while the latter excludes the weakly magnetized disk. In the following, we call the region identified by the two conditions as the ‘dissipative jet sheath,’ or the ‘candidate corona.’
@@ 3.2.2 Azimuthal field strength in the reconnection layers
The figure shows that the azimuthal field in the northern hemisphere is preferentially negative (the opposite holds in the southern hemisphere), although some dissipation regions in the jet sheath have mostly positive $B_{\rm out}$ (see third and fourth panels). During flaring events, magnetic loops originating in the southern hemisphere (and so, preferentially having $B_{\rm out}>0$) may be launched into the jet sheath of the upper hemisphere, as illustrated by the third and fourth panels in Fig. 5. We also point out that at the equatorial current sheet appearing during flux eruptions, all field components (i.e., both in-plane fields as well as $B_{\rm out}$) flip their polarity (Ripperda et al., 2020, 2022), i.e., here reconnection proceeds in the regime of vanishing guide field.
@@
If the in-plane field reverses polarity across a current sheet, while the out-of-plane component does not, then one could interpret the out-of-plane component—along the direction of the electric current—as a guide field. While $B_{\rm out}$ does not always act as a guide field in the dissipative jet sheath, it could locally act as a guide field at certain times, as its polarity often stays constant across the current sheets in the jet sheath (e.g., see first and second panels of Fig. 5). [^3] This is true in our two-dimensional simulation and may change in three-dimensional simulations where physical quantities will not be invariant in the $\phi$ direction. On the contrary, as already anticipated above, across the equatorial current sheet all field components reverse in polarity, so the out-of-plane azimuthal field cannot be interpreted as a guide field there. The properties of particle energization, reconnection plasmoids, and their bulk motions depend on the strength of the out-of-plane guide field, as demonstrated by local kinetic studies of reconnection (e.g., Ball et al., 2018; Gupta et al., 2024). Large guide-field reconnecting current sheets in the jet sheath might also be sites of proton acceleration that can produce high-energy neutrinos (Fiorillo et al., 2023).
@@ 3.3 Power in the jet sheath
We find that along the jet sheath, $\dot{E}^{\rm(EM)}(r)$ decreases with radius as $\propto r^{-0.2}$ from $2-10\,R_{\rm g}$ (demonstrating magnetic dissipation) followed by an increase with radius (shown in Fig. 14 until $\sim 40\,R_{\rm g}$): note that an increasing $\dot{E}^{\rm(EM)}(r)$ at larger radii does not mean there is no dissipation there, instead, it could be due to the additional contributions from disk winds (as discussed in Qian et al., 2018). For the parameters typical of Cygnus X-1 (mentioned below in §3.5), we find that up to $\sim 10^{38}\,{\rm erg\,s^{-1}}$ may be dissipated along the jet sheath from $2-10\,R_{\rm g}$.
@@ 3.4 Bulk motions in the dissipative jet sheath
Figure 9: Left panel: Time- and $\theta$-averaged bulk energy spectra, calculated in the local comoving frame between 5 $R_{\rm g}$ (blue) and 40 $R_{\rm g}$ (red), in bins of 5 $R_{\rm g}$. The black dotted and dashed curves denote 100 keV and 200 keV Maxwellian distributions, respectively. Right panel: Comoving-frame mean bulk energy (solid line) and its standard deviation (shaded band) as a function of radius within the dissipative jet sheath. All quantities are measured in the fluid comoving frame within the dissipative jet sheath in the upper hemisphere and averaged over the quasi-steady state.
In the left panel of Fig. 8, we show the radial ($\hat{r}$; red), polar ($\hat{\theta}$; green), and azimuthal ($\hat{\phi}$; blue) components of the dimensionless Eulerian bulk 4-velocity ($\Gamma\boldsymbol{\beta}$) as a function of radius. All quantities are averaged over the polar angle within the dissipative jet sheath and in time during the quasi-steady state. The colored shaded band around each curve denotes the corresponding standard deviation. We find that the fluid bulk motions are dominated by the radial component ($\Gamma\beta_{\rm r}$), with $\langle\Gamma\beta_{\rm r}\rangle\pm\sigma_{\Gamma\beta_{\rm r}}$ that increases from $1.0\pm 0.4$ at $r=5\,R_{\rm g}$ until it saturates at $1.3\pm 0.6$ for $r\gtrsim 30\,R_{\rm g}$. The polar component decreases from $\langle\Gamma\beta_{\rm\theta}\rangle\pm\sigma_{\Gamma\beta_{\rm\theta}}=-0.12\pm 0.11$ at $r=5\,R_{\rm g}$ to $-0.24\pm 0.11$ at $r=40\,R_{\rm g}$. The azimuthal component also decreases with radius, from $\langle\Gamma\beta_{\rm\phi}\rangle\pm\sigma_{\Gamma\beta_{\rm\phi}}=0.43\pm 0.28$ at $r=5\,R_{\rm g}$ to $0.16\pm 0.15$ at $r=40\,R_{\rm g}$. The mean bulk motion of the dissipative jet sheath, as quantified by the left panel of Fig. 8, is crucial in determining (i) the energy density and angular distribution of disk photons getting reprocessed in the corona (e.g., Zdziarski et al. 1999), and (ii) the fraction of Comptonized X-rays that return to the disk and give rise to the reflected component (e.g., Beloborodov 1999). Our result of a radially-accelerating coronal region is in agreement with the MHD models of Merloni & Fabian (2002) and with the empirical model of King et al. (2017).
@@
for $a_{\bullet}=0.9375$. At the jet boundary ($\theta\sim\theta_{\rm pk}$), far away from the black hole, we have $\Omega r\sin\theta_{\rm pk}\gg 1$, which implies $E\gg B_{\rm p}$. On the other hand, at most $E\simeq B_{\rm out}$. This implies that the drift Lorentz factor $\Gamma_{\rm d}$ can be calculated as
@@
So far we have characterized the mean bulk motions. Within the paradigm of cold-chain Comptonization, the properties of the Comptonized hard nonthermal X-ray emission are determined by the stochasticity in the plasmoid motions. To this end, we compute the polar- and time-averaged bulk energy spectra within the dissipative jet sheath, in the frame moving with the local mean velocity (as quantified above). The histogram of bulk motions measured in this frame is shown by the solid curves in the left panel of Fig. 9 as a function of radial distance, from $5\,R_{\rm g}$ (blue) to $40\,R_{\rm g}$ (red) in increments of $5\,R_{\rm g}$. The distribution of stochastic bulk motions resembles a Maxwellian with an ‘effective bulk temperature’ of $\sim 200$ keV at $5\,R_{\rm g}$, decreasing down to $\sim 100$ keV at $40\,R_{\rm g}$ (see dashed and dotted black curves). In the right panel, we show the average bulk Lorentz factor $\langle\Gamma^{\prime}-1\rangle$ and its standard deviation $\sigma_{\Gamma^{\prime}-1}$ (as shown by the dark grey band), both measured in the frame moving with the local mean velocity. We find that $\langle\Gamma^{\prime}-1\rangle\pm\sigma_{\Gamma^{\prime}-1}$ decrease from $0.40\pm 0.14$ at $r=10\,R_{\rm g}$ to $0.3\pm 0.1$ at $r=40\,R_{\rm g}$.
In the traditional picture of thermal Comptonization, hot electrons with a temperature of $\sim 100$ keV are required in order to explain the observed hard nonthermal X-ray spectra. In our model, it is the stochasticity in the bulk motions of dissipative structures within the jet sheath—remarkably, having an effective bulk temperature of $\sim 100$ keV—that allows to Compton upscatter soft photons to a hard nonthermal X-ray distribution. In previous works, we have employed Monte Carlo Radiative Transfer calculations to show that bulk motions with an effective bulk temperature of $\sim 100$ keV are capable to explain the observed hard nonthermal X-ray spectra through the cold-chain Comptonization process (SSB21, SSB23). In local simulations of relativistic reconnection, the stochasticity of bulk motions is a natural by-product of the plasmoid chain dynamics (e.g., SSB21, SSB23). Here, our choice of $\sigma_{\rm h}>0.15$ and $\beta>0.1$ to identify dissipative regions in the jet sheath is tailored towards capturing reconnection layers and plasmoids. Still, stochastic motions may be contributed by other sources, e.g., the waves/vortices seen at the jet sheath (check panels [a,b] of Fig. 2 for the presence of nonlinear waves along the jet sheath).
@@ 3.5 Optical depth
$\tau_{\rm e^{-}p}\sim\left(\frac{w}{0.1\,r}\right)\,\frac{2\ell_{\rm B}/\sigma_{\rm c}}{m_{\rm p}/m_{\rm e}}.$ (14)
@@
for a comoving field $b$ in simulation units. In Eq. (14), we assume that the width of the current sheets in the jet sheath to be $w\sim 0.1\,r$. In the case of plasmoid chain Comptonization, Beloborodov (2017) adopted a similar definition, where $0.1$ was the aspect ratio of the reconnection layer, set by the reconnection rate. We remark that Eq. (14) is a reasonable estimate of optical depth only if the mass density in the sheath (in $\sigma_{\rm c}$) is not dominated by the density floors of the GRRMHD algorithm.
@@
Figure 10: Polar- and time-averaged radial profiles of the comoving magnetic energy density (red), the magnetic compactness (green), the optical depth of electron-positron pairs (blue) and of electron-proton plasma (brown) in the dissipative jet sheath, as obtained from our simulation assuming parameters typical of Cygnus X-1 (see §3.5 for more details).
where we adopt $\xi=0.2$ as the efficiency of electromagnetic dissipation (see §3.3 and Appendix C). In Fig. 10, compute the polar-averaged and time-averaged profiles of $\ell_{\rm B}$, $\tau_{\rm e^{-}p}$ and $\tau_{\rm e^{\pm}}$ as a function of radial distance from the black hole, considering only regions belonging to the dissipative jet sheath (i.e., $\beta>0.1$ and $\sigma_{\rm h}>0.15$). We find that, for a (putative) corona at $2\lesssim r/R_{\rm g}\lesssim 20$, the electron-proton optical depth is $\tau_{\rm e^{-}p}\sim 10^{-3}-10^{-2}$, while the self-regulated pair optical depth is $\tau_{\rm e^{\pm}}\sim 0.01-0.1$. Our estimate of $\tau_{\rm e^{-}p}$ is affected by the threshold $\sigma_{\rm h}>0.15$ used to identify the dissipative jet sheath; a less restrictive threshold will likely give higher $\tau_{\rm e^{-}p}$. We note that future global GRRMHD simulations with self-consistent radiative physics are needed for a proper estimate of the optical depth in the dissipative jet sheath.

### v2

@@ 3 GRRMHD simulation results
Figure 3: Realizations of plasmoid-mediated reconnection layers identified in the jet sheath by the conditions $\sigma_{\rm h}>0.15$ and $\beta>0.1$ at various times (in different rows). Panels [a-d] show $\sigma_{\rm h}$; panels [e-h] show $\beta$; panels [i-l] show the current density. The bulk energy per particle in this region (i.e., where $\sigma_{\rm h}>0.15$ and $\beta>0.1$) is shown in panels [m-p]. In all panels, the black hole is centered at $[x,z]=[0,0]$, the red dashed curve is the ergosphere, and the red solid curve is the inner light surface. Click here for a YouTube video of some of these quantities.
@@ 3.1 General properties of the jet sheath
In the top panel of Fig. 2, we present two-dimensional snapshots of the plasma density, the dimensionless temperature ($p/\rho c^{2}$), and the bulk energy per unit rest mass energy ($\Gamma-1$) at time $Tc/R_{\rm g}=2870$. One can identify the jet sheath as the region where the dimensionless temperature is $p/\rho c^{2}\sim 0.1$ and the bulk energy is $\Gamma-1\sim 1$. In the jet sheath, the flow transitions from being ultra-relativistically fast and hot in the jet to non-relativistic and cold in the disk wind. The plasma mass density is $\rho\sim 0.1$ (in code units) at the jet-disk interface. In the following sections, we present a different criterion—based on $\sigma_{\rm h}$ and $\beta$—to identify the regions with reconnection-like features in the jet sheath.
@@ 3.2.1 Location and geometry
Note that the presence of a current sheet does not guarantee fast dissipation. However, a current sheet showing plasmoid chains is undergoing active reconnection, which implies fast dissipation. Most of the time, long current sheets in the jet sheath are seen to harbor chains of plasmoids, which appear as prominent regions with $\beta\gtrsim 1$. Most of the plasmoids residing in/near the sheath at a given time were born in the sheath itself. However, there are brief moments (of duration $\sim 10\,R_{\rm g}/c$) when plasmoids are not formed in situ in the sheath: these are the times immediately following a flux eruption. During those moments, the region adjacent to the sheath (toward the disk, with $\sigma_{\rm h}<0.15$) may still harbor large plasmoids ($\sim 5\,R_{\rm g}/c$), which were created at the base of the jet during the prior flux eruption. These plasmoids can grow to large sizes (${\cal O}(10)\,R_{\rm g}$) while propagating outwards. Some examples appear in the top panel of Fig. 2 as the closed loops of magnetic field lines (e.g., at $[x/R_{\rm g},z/R_{\rm g}]=[16,7],[24,22]$). Numerical simulations have revealed that some of these large plasmoids move along a spiral trajectory around the jet (Nathanail et al., 2020; Ripperda et al., 2022; El Mellah et al., 2023).
Given the copious presence of plasmoids and reconnecting layers in the jet sheath, we regard this interface as an important site of energy dissipation. As such, it may be a potential candidate for the Comptonizing corona. More quantitatively, we identify the reconnection-powered dissipation regions in the jet sheath by the following combination: plasma beta $\beta>1$ and hot magnetization $\sigma_{\rm h}>0.15$. The former threshold excludes the jet core (which has lower plasma beta, and is unphysically affected by numerical density and pressure floors), while the latter excludes the weakly magnetized disk. In the following, we call the region identified by the two conditions the ‘dissipative jet sheath,’ or the ‘candidate corona.’
@@ 3.2.2 Azimuthal field strength in the reconnection layers
The figure shows that the azimuthal field in the northern hemisphere is preferentially negative (the opposite holds in the southern hemisphere), although some dissipation regions in the jet sheath have mostly positive $B_{\rm out}$ (see third and fourth panels). During flaring events, magnetic loops originating in the southern hemisphere (and thus preferentially having $B_{\rm out}>0$) may be launched into the jet sheath of the upper hemisphere, as illustrated by the third and fourth panels in Fig. 5. We also point out that at the equatorial current sheet appearing during flux eruptions, all field components (i.e., both in-plane fields and $B_{\rm out}$) flip their polarity (Ripperda et al., 2020, 2022), i.e., here reconnection proceeds in the regime of vanishing guide field.
@@
If the in-plane field reverses polarity across a current sheet, while the out-of-plane component does not, then one could interpret the out-of-plane component—along the direction of the electric current—as a guide field. While $B_{\rm out}$ does not always act as a guide field in the dissipative jet sheath, it could locally act as a guide field at certain times, as its polarity often stays constant across the current sheets in the jet sheath (e.g., see first and second panels of Fig. 5). [^3] This is true in our two-dimensional simulation and may change in three-dimensional simulations where physical quantities will not be invariant in the $\phi$ direction. In contrast, as already anticipated above, across the equatorial current sheet all field components reverse in polarity, so the out-of-plane azimuthal field cannot be interpreted as a guide field there. The properties of particle energization, reconnection plasmoids, and their bulk motions depend on the strength of the out-of-plane guide field, as demonstrated by local kinetic studies of reconnection (e.g., Ball et al., 2018; Gupta et al., 2024). Current sheets that undergo reconnection in the presence of large guide fields in the jet sheath might also be sites of proton acceleration that can produce high-energy neutrinos (Fiorillo et al., 2023).
@@ 3.3 Power in the jet sheath
We find that along the jet sheath, $\dot{E}^{\rm(EM)}(r)$ decreases with radius as $\propto r^{-0.2}$ from $2-10\,R_{\rm g}$ (demonstrating magnetic dissipation) then increases with radius (shown in Fig. 14 until $\sim 40\,R_{\rm g}$): note that an increasing $\dot{E}^{\rm(EM)}(r)$ at larger radii does not mean there is no dissipation there; instead, it could be due to the additional contributions from disk winds (as discussed in Qian et al., 2018). For the parameters typical of Cygnus X-1 (mentioned below in §3.5), we find that up to $\sim 10^{38}\,{\rm erg\,s^{-1}}$ may be dissipated along the jet sheath from $2-10\,R_{\rm g}$.
@@ 3.4 Bulk motions in the dissipative jet sheath
Figure 9: Left panel: time- and $\theta$-averaged bulk energy spectra, calculated in the local comoving frame between 5 $R_{\rm g}$ (blue) and 40 $R_{\rm g}$ (red), in bins of 5 $R_{\rm g}$. The black dotted and dashed curves denote 100 keV and 200 keV Maxwellian distributions, respectively. Right panel: mean bulk energy in the comoving frame (solid line) and its standard deviation (shaded band) as a function of radius within the dissipative jet sheath. All quantities are measured in the frame comoving with the fluid within the dissipative jet sheath in the upper hemisphere and averaged over the quasi-steady state.
In the left panel of Fig. 8, we show the radial ($\hat{r}$; red), polar ($\hat{\theta}$; green), and azimuthal ($\hat{\phi}$; blue) components of the dimensionless Eulerian bulk 4-velocity ($\Gamma\boldsymbol{\beta}$) as a function of radius. All quantities are averaged over the polar angle within the dissipative jet sheath and in time during the quasi-steady state. The colored shaded band around each curve denotes the corresponding standard deviation. We find that the fluid bulk motions are dominated by the radial component ($\Gamma\beta_{\rm r}$), with $\langle\Gamma\beta_{\rm r}\rangle\pm\sigma_{\Gamma\beta_{\rm r}}$, which increases from $1.0\pm 0.4$ at $r=5\,R_{\rm g}$ until it saturates at $1.3\pm 0.6$ for $r\gtrsim 30\,R_{\rm g}$. The polar component decreases from $\langle\Gamma\beta_{\rm\theta}\rangle\pm\sigma_{\Gamma\beta_{\rm\theta}}=-0.12\pm 0.11$ at $r=5\,R_{\rm g}$ to $-0.24\pm 0.11$ at $r=40\,R_{\rm g}$. The azimuthal component also decreases with radius, from $\langle\Gamma\beta_{\rm\phi}\rangle\pm\sigma_{\Gamma\beta_{\rm\phi}}=0.43\pm 0.28$ at $r=5\,R_{\rm g}$ to $0.16\pm 0.15$ at $r=40\,R_{\rm g}$. The mean bulk motion of the dissipative jet sheath, as quantified by the left panel of Fig. 8, is crucial in determining (i) the energy density and angular distribution of disk photons getting reprocessed in the corona (e.g., Zdziarski et al. 1999), and (ii) the fraction of Comptonized X-rays that return to the disk and give rise to the reflected component (e.g., Beloborodov 1999). Our result of a radially-accelerating coronal region is in agreement with the MHD models of Merloni & Fabian (2002) and with the empirical model of King et al. (2017).
@@
for $a_{\bullet}=0.9375$. At the jet boundary ($\theta\sim\theta_{\rm pk}$), far away from the black hole, we have $\Omega r\sin\theta_{\rm pk}\gg 1$, which implies $E\gg B_{\rm p}$. On the other hand, $E\simeq B_{\rm out}$ at most. This implies that the drift Lorentz factor $\Gamma_{\rm d}$ can be calculated as
@@
So far we have characterized the mean bulk motions. Within the paradigm of cold-chain Comptonization, the properties of the Comptonized hard nonthermal X-ray emission are determined by the stochasticity in the plasmoid motions. To this end, we compute the polar- and time-averaged bulk energy spectra within the dissipative jet sheath, in the frame moving with the local mean velocity (as quantified above). The histogram of bulk motions measured in this frame is shown by the solid curves in the left panel of Fig. 9 as a function of radial distance, from $5\,R_{\rm g}$ (blue) to $40\,R_{\rm g}$ (red) in increments of $5\,R_{\rm g}$. The distribution of stochastic bulk motions resembles a Maxwellian with an ‘effective bulk temperature’ of $\sim 200$ keV at $5\,R_{\rm g}$, decreasing down to $\sim 100$ keV at $40\,R_{\rm g}$ (see dashed and dotted black curves). In the right panel, we show the average bulk Lorentz factor $\langle\Gamma^{\prime}-1\rangle$ and its standard deviation $\sigma_{\Gamma^{\prime}-1}$ (as shown by the dark grey band), both measured in the frame moving with the local mean velocity. We find that $\langle\Gamma^{\prime}-1\rangle\pm\sigma_{\Gamma^{\prime}-1}$ decreases from $0.40\pm 0.14$ at $r=10\,R_{\rm g}$ to $0.3\pm 0.1$ at $r=40\,R_{\rm g}$.
In the traditional picture of thermal Comptonization, hot electrons with a temperature of $\sim 100$ keV are required in order to explain the observed hard nonthermal X-ray spectra. In our model, it is the stochasticity in the bulk motions of dissipative structures within the jet sheath—remarkably, having an effective bulk temperature of $\sim 100$ keV—that allows soft photons to be Compton scattered to a hard nonthermal X-ray distribution. In previous works, we have employed Monte Carlo radiative transfer calculations to show that bulk motions with an effective bulk temperature of $\sim 100$ keV are capable of explaining the observed hard nonthermal X-ray spectra through the cold-chain Comptonization process (SSB21, SSB23). In local simulations of relativistic reconnection, the stochasticity of bulk motions is a natural by-product of the plasmoid chain dynamics (e.g., SSB21, SSB23). Here, our choice of $\sigma_{\rm h}>0.15$ and $\beta>0.1$ to identify dissipative regions in the jet sheath is tailored toward capturing reconnection layers and plasmoids. Still, stochastic motions may be contributed by other sources, e.g., the waves/vortices seen at the jet sheath (check panels [a,b] of Fig. 2 for the presence of nonlinear waves along the jet sheath).
@@ 3.5 Optical depth
$\tau_{\rm e^{-}p}\sim\left(\frac{w}{0.1\,r}\right)\,\frac{2\ell_{\rm B}/\sigma_{\rm c}}{m_{\rm p}/m_{\rm e}},$ (14)
@@
for a comoving field $b$ in simulation units. In Eq. (14), we assume the width of the current sheets in the jet sheath to be $w\sim 0.1\,r$. In the case of plasmoid chain Comptonization, Beloborodov (2017) adopted a similar definition, where $0.1$ was the aspect ratio of the reconnection layer, set by the reconnection rate. We remark that Eq. (14) is a reasonable estimate of optical depth only if the mass density in the sheath (in $\sigma_{\rm c}$) is not dominated by the density floors of the GRRMHD algorithm.
@@
Figure 10: Polar- and time-averaged radial profiles of the comoving magnetic energy density (red), the magnetic compactness (green), the optical depth of electron-positron pairs (blue) and of electron-proton plasma (brown) in the dissipative jet sheath, as obtained from our simulation scaled to the parameters typical of Cygnus X-1 (see §3.5 for more details).
where we adopt $\xi=0.2$ as the efficiency of electromagnetic dissipation (see §3.3 and Appendix C). Fig. 10 shows the polar-averaged and time-averaged profiles of $\ell_{\rm B}$, $\tau_{\rm e^{-}p}$ and $\tau_{\rm e^{\pm}}$ as a function of radial distance from the black hole, considering only regions belonging to the dissipative jet sheath (i.e. where $\beta>0.1$ and $\sigma_{\rm h}>0.15$). We find that, for a (putative) corona at $2\lesssim r/R_{\rm g}\lesssim 20$, the electron-proton optical depth is $\tau_{\rm e^{-}p}\sim 10^{-3}-10^{-2}$, while the self-regulated pair optical depth is $\tau_{\rm e^{\pm}}\sim 0.01-0.1$. Our estimate of $\tau_{\rm e^{-}p}$ is affected by the threshold $\sigma_{\rm h}>0.15$ used to identify the dissipative jet sheath; a less restrictive threshold will likely give higher $\tau_{\rm e^{-}p}$. We note that future global GRRMHD simulations with self-consistent radiative physics are needed for a proper estimate of the optical depth in the dissipative jet sheath.